the perimeter of a rectangle is 48 inches.The length is 6 inches greater than the width.Define the variables,and write equaations to represent this situation.solve the system by substitution
Answer by tsewaitung(11)

let the length be x and the width be y.
from the question, we have
2(x+y)=48
x=y+6
by substituting the second equaation into the first equaation, we have
2[(y+6)+y)=48
2[y+6+y]=48
2[2y+6]=48
4[y+3]=48
y+3=12
y=9
by substituting y=9 into the second equaation, we have
x=9+6
x=15
the length of the rectangle is 15 inches and the width of the rectangle is 9 inches.